The mass of a cork is 2g.What is the fraction o

f the cork is immersed when it floats in water

1
Expert's answer
2021-05-18T11:10:06-0400

"m = 2*10^{-3}kg"

"\rho= 1*10^3\frac{kg}{m^3}"

"g\approx 10\frac{m}{s^2}"

"F_a = \rho gV_{im}=10^4V_{im}"

"F_g= mg=2*10^{-2}"

"F_g =F_a"

"10^4V_{im}=2*10^{-2}"

"V_{im}=2*10^{-6}m^3\text{The volume of a body immersed in a liquid}"

"\text{Answer: }2*10^{-6}\ m^3"